OpenStudy (anonymous):

Question 2. 2. Solve, using linear combination. 4x + y = 5 3x + y = 3 (Points : 1) (0, –3) (0, 5) (2, –3) (–3, 7)

OpenStudy (mathmale):

If: 4x + y = 5 3x + y = 3 Multiply the 2nd equation all the way through (evey term) by -1. Then combine that result with 4x + y = 5 Note: y will "disappear." Find x.

OpenStudy (anonymous):

so the answer (-3,7)

OpenStudy (mathmale):

If you'd like to check your result, substitute the x- and y-coordinates of that result into either of the given equations. Is the equation then true or false? To be thorough, check your result in the other equation as well.

OpenStudy (anonymous):

u lost me

OpenStudy (mathmale):

You have a choice: You could take the coordinates of each of the given answer choices and substitute its x-coordinate for x in 4x + y = 5 and its y-coordinate for y in that same equation. Either the resulting equation is TRUE or it is not. Or, you could solve the system of linear equations yourself and then compare your result with the four answer choices. Which will it be?